OpenStudy (anonymous):

Tell me the answer to this equation plz: will award medal 17^2-x=20

OpenStudy (mwhit2101):

17^2 is 289 so your equation is: 289-x=20 Subtract 289 from both sides -X=-269 There is an understood 1 in x so divide both sides by 1 -x/-1 and -269/-1 X=269 Answer ^^^

OpenStudy (anonymous):

17*17 = 289 289 -20 = ?

OpenStudy (anonymous):

Thank you both

OpenStudy (anonymous):

peanut butter